### v1

# MoLoRec: A Generalizable and Efficient Framework for LLM-Based Recommendation

Abstract.
Large Language Models (LLMs) have achieved remarkable success in recent years, owing to their impressive generalization capabilities and rich world knowledge. To capitalize on the potential of using LLMs as recommender systems, mainstream approaches typically focus on two paradigms. The first paradigm designs multi-domain or multi-task instruction data for generalizable recommendation, so as to align LLMs with general recommendation areas and deal with cold-start recommendation. The second paradigm enhances domain-specific recommendation tasks with parameter-efficient fine-tuning techniques, in order to improve models under the warm recommendation scenarios. While most previous works treat these two paradigms separately, we argue that they have complementary advantages, and combining them together would be helpful.
To that end, in this paper, we propose a generalizable and efficient LLM-based recommendation framework MoLoRec. Our approach starts by parameter-efficient fine-tuning a domain-general module with general recommendation instruction data, to align LLM with recommendation knowledge. Then, given users’ behavior of a specific domain, we construct a domain-specific instruction dataset and apply efficient fine-tuning to the pre-trained LLM. After that, we provide approaches to integrate the above domain-general part and domain-specific part with parameters mixture. Please note that, MoLoRec is efficient with plug and play, as the domain-general module is trained only once, and any domain-specific plug-in can be efficiently merged with only domain-specific fine-tuning. Extensive experiments on multiple datasets under both warm and cold-start recommendation scenarios validate the effectiveness and generality of the proposed MoLoRec. Codes are available at https://anonymous.4open.science/r/MoLoRec.
1. Introduction
Large Language Models (LLMs) have demonstrated significant success across diverse fields (Zhao et al., 2023), driven by their emergent capabilities (Dong et al., 2024; Huang and Chang, 2023) such as world knowledge, language understanding, and complex reasoning. Recently, LLMs have introduced transformative advancements to recommendation tasks. Notably, LLMs have shown potential in capturing nuanced item semantics (Xi et al., 2024), understanding diverse user interests (Deng et al., 2022), and unifying various recommendation tasks (Geng et al., 2022). These advancements highlight the promise of utilizing LLMs as recommender systems, positioning LLM-based recommendations as a compelling area for further exploration.
In the field of LLM-based recommendation research, the emergence of ChatGPT and its remarkable reasoning capabilities have catalyzed early studies (Dai et al., 2023; Sanner et al., 2023; Wang et al., 2023). These works focus on the zero-shot/few-shot recommendation potential of LLMs through in-context learning (Dong et al., 2024). However, the intrinsic gap between the pre-training general text corpus of LLMs and the requirements of recommendation tasks results in suboptimal performance when relying solely on in-context learning. Consequently, the key to developing an effective LLM-based recommender system lies in bridging this gap, enabling the model to truly ”understand” how to recommend. To address this challenge, researchers have proposed a variety of approaches. We classify them into two paradigms, each tackling the problem from a distinct perspective. As shown in Figure 1(a), the first one is summarized as the breadth-oriented paradigm. These works integrate multi-domain (Tang et al., 2024) or multi-task (Geng et al., 2022; Zhang et al., 2024; Cui et al., 2022) recommendation data to construct extensive recommendation world knowledge, paving the way for developing a generalizable LLM-based recommender. The key focus of this paradigm is the integration of multi-source data to build instruction-tuning datasets (Peng et al., 2024; Jin et al., 2023) and the design of instruction templates (Zhang et al., 2024; Geng et al., 2022) tailored to various tasks. The second paradigm is termed the depth-oriented paradigm, illustrated in Figure 1(b). This line of research seeks to enable LLMs to deeply comprehend recommendation tasks within specific domains. Key areas of focus include: the in-depth extraction of domain-specific recommendation knowledge, such as collaborative filtering information (Lin et al., 2024b; Kim et al., 2024a; Liao et al., 2024; Ren et al., 2024; Kong et al., 2024), and the development of efficient and effective alignment methods between LLMs and recommendation tasks. Specifically, compared with enormous parameters in LLMs, downstream tasks do not have sufficient data for tuning all parameters. Therefore, parameter-efficient fine-tuning methods become optimal for applying LLMs, in which lightweight Low-Rank Adapter (LoRA) is one representative work (Hu et al., 2022). By borrowing ideas of LLMs, these methods include leveraging (Bao et al., 2023) or enhancing (Kong et al., 2024) LoRA fine-tuning techniques and designing data-efficient fine-tuning strategies (Lin et al., 2024c).
These works make significant advancements in recommendation research. Nevertheless, we argue that these two paradigms have complementary advantages. The first paradigm masters general recommendation knowledge and can be well generalized to various recommendation scenarios. The second paradigm learns each user’s unique preference and is suitable for the warm-start recommendation of a specific domain. However, the specific recommendation domain performance could not be easily transferred to other domains. In fact, both generalizable recommendation knowledge and efficient domain-specific understanding are essential for recommender systems. Relying solely on one aspect risks falling short in addressing the diverse challenges in real-world recommendation scenarios. The breadth-oriented paradigm may underperform in specific domains. Conversely, the depth-oriented paradigm struggles with distribution shifts between training and test data. It faces challenges when new users or items appear or when training data are sparse.
To this end, we investigate how to integrate the advantages of both paradigms to simultaneously enhance the model’s generalization ability and domain-specific performance. The task creates significant obstacles: (1) Efficiency. Integrating two paradigms may introduce model complexity, and finding an efficient integrating method without excessive computational overhead is a critical challenge. (2) Generalizability. We need to preserve the model’s generalization ability to a large extent, enabling it to quickly scale to new domains, new items, and other new recommendation scenarios. In this paper, we propose a generalizable and efficient recommendation framework named Mixture-of-LoRA Recommendation Framework (MoLoRec). MoLoRec fuses both general recommendation knowledge and domain-specific knowledge with three key stages. Firstly, to align LLM with any recommendation task, MoLoRec constructs a general recommendation instruction dataset from multiple recommendation domains, and fine-tunes LLM to get a domain-general LoRA module. Secondly, to tailor the framework for specific domains, MoLoRec constructs domain-specific instruction datasets derived from the specific domain, and fine-tunes LLM to get a domain-specific LoRA module. After that, MoLoRec performs a highly efficient and effective linear arithmetic operation to merge these LoRA adapters within the weight space, allowing MoLoRec to maintain strong recommendation performance across both specific domains and out-of-distribution scenarios. To further enhance the merging process of LoRA adapters, we also introduce an adaptive LoRA weight merging method guided by entropy minimization during test time. Importantly, our framework is designed for ease of use, allowing for a plug-and-play integration where the domain-general module is trained once, and domain-specific adaptations are incorporated through minimal fine-tuning. Finally, extensive experiments conducted on various datasets demonstrate the effectiveness and versatility of the framework in both warm recommendation scenarios and challenging cold-start scenarios, highlighting its potential for broad application.
2. Preliminary
In this section, we introduce key concepts underpinning our methodology. First, we cover the task formulation and instruction tuning for LLM-based recommender models. Next, we highlight the use of LoRA for parameter-efficient fine-tuning of LLMs.
2.1. LLM-Based Recommender Models
Task Formulation. We mainly focus on the sequential recommendation task, which holds significant practical importance. Let and represent the sets of users and items, respectively. The historical interaction sequence of a user is denoted as , arranged in chronological order, where and . The goal is to predict this user’s next liked item based on the historical interactions.
Instruction Tuning for LLM-Based Recommendation. For LLM-based recommendation, instruction tuning (Wei et al., 2022) is the key step to bridge the gap between the next-word prediction objective of LLMs and the recommendation task (Bao et al., 2023; Kong et al., 2024; Lin et al., 2024c). Formally, instruction tuning involves fine-tuning LLMs using training data organized as explicit instruction pairs . Here, represents a detailed textual instruction that encapsulates the interaction sequences and the recommendation task, while corresponds to the textual description of the predicted item . The instruction fine-tuning process is guided by minimizing the following autoregressive loss function:
where denotes the -th token of , is the token sequence preceding , and is the LLM’s parameters.
2.2. Low-Rank Adaptation (LoRA)
In traditional fine-tuning as described in Eqn. (1), updating all parameters makes the process highly computationally intensive, particularly for LLMs. To address this issue, parameter-efficient methods are designed to fine-tune LLMs while updating only a small subset of parameters. Low-Rank Adaptation (LoRA) (Hu et al., 2022) is the mainstream approach. LoRA addresses this issue by introducing low-rank matrices that are trained alongside the frozen original model weights. This allows the model to adapt to specific tasks by learning a small number of additional parameters, without requiring modifications to the entire model.
Specifically, for any pre-trained weight matrics in the transformer block of the LLM, which takes an input and output . LoRA modifies to:
where , are the low-rank projection matrices. Notably, the rank , ensuring that the number of parameters introduced by is significantly fewer than those of , as . During fine-tuning, only and are updated, while remains fixed. In a similar way, LoRA adapter is generally applicable to any LLM layer desired for updating. The training objective of LoRA fine-tuning can be formulated as:
Here, is the parameters of the pre-trained LLM. denotes the set of parameters of LoRA fine-tuning, and represents the number of LoRA modules.
3. Methodology
In this section, we propose MoLoRec, a generalizable, effective, and efficient LLM-based recommendation framework. As shown in Figure 2, MoLoRec operates through three key stages. Firstly, we align the LLM with any recommendation task and train a domain-general LoRA module (Section 3.1). Secondly, to adapt the framework to specific domains, we fine-tunes the LLM to obtain domain-specific LoRA modules (Section 3.2). Subsequently, MoLoRec performs an efficient linear arithmetic operation to merge these LoRA adapters within the weight space, enabling it to maintain strong recommendation performance across both specific domains and out-of-distribution scenarios (Section 3.3). Notably, our framework is designed for ease of use, offering plug-and-play integration, where the domain-general module is trained once, and domain-specific adaptations are incorporated with minimal fine-tuning. MoLoRec is capable of significantly enhancing the recommendation performance and generalization across various domains at a relatively low cost.
3.1. Training Generalizable Base Recommender
Pre-trained LLMs often exhibit suboptimal performance when directly applied to recommendation tasks (Bao et al., 2023; Lin et al., 2024a). This limitation arises because their pre-training on general-purpose datasets fails to capture the specialized knowledge required for understanding user preferences, behaviors, and the contextual nuances essential to building effective recommender systems. To bridge this gap, in this subsection, we align the pre-trained LLM with the recommendation task, as shown in Figure 2(a). Specifically, we first construct large-scale instruction data by mixing users’ behaviors from multiple domains. The pre-trained LLM is then fine-tuned on this dataset using the LoRA technique, equipping it with general knowledge to tackle recommendation tasks.
Instruction Data Construction. Given recommendation domains (i.e., , ,…, ), denotes , , and as the user set, item set and user interaction sequence set of domain , respectively. To provide general user modeling and recommendation knowledge, we combine data from all domains and design instruction templates to convert them into a text format. Note that the choice of recommendation domains and instruction templates can be arbitrary. As illustrated in Figure 2(a), we transform the recommendation data into instruction data , where x and y denote the instruction input and output, respectively. The instruction input includes the task description (which explains the recommendation task), the user’s historical interactions, and the item candidate set, all in natural language. Here, the items are represented by their titles. The candidate set consists of one ground-truth item and some randomly selected negative samples. The instruction output is designed to rank the user’s next most likely products.
Tuning Domain-General LoRA Module. Given the instruction data , we apply LoRA fine-tuning to adapt the pre-trained LLM for general recommendation tasks. The pre-trained model parameters are kept frozen, while trainable low-rank decomposition matrices are introduced into each layer of the Transformer architecture, enabling efficient and lightweight tuning. Formally,
where is the parameters of the pre-trained LLM, and denotes the set of parameters of LoRA fine-tuning. By undergoing this fine-tuning step, is now enriched with extensive general knowledge relevant to the field of recommendations.
3.2. Training Domain-Specific Plugin
Each recommendation domain exhibits unique user behavior patterns, making the acquisition of domain-specific knowledge essential for delivering accurate recommendations. For instance, in the clothing recommendation domain, user behavior is predominantly driven by style preferences. In contrast, when purchasing electronic products, users tend to prioritize compatibility with their existing products. To address these domain-specific needs, we construct an instruction dataset tailored to the specific new recommendation domain and apply LoRA fine-tuning to the pre-trained LLM, resulting in the LoRA modules enhanced with domain-specific expertise. As shown in Figure 2(b), the template for constructing the instruction dataset and LoRA fine-tuning method are similar to those outlined in Section 3.1. We formally define the process of training the domain-specific LoRA module as:
3.3. Mixture-of-LoRA for Plug-and-Play.
After instruction fine-tuning in stages 1 and 2, we obtained a domain-general model that captures general recommendation knowledge and a domain-specific model that incorporates domain-specific insights. In this subsection, we propose integrating these two parts to improve recommendation accuracy and enhance generalization capabilities simultaneously. Natural questions arise: could this goal be achieved by applying traditional ensemble learning methods that combine the outputs of multiple models? Or could we integrate general and domain-specific knowledge by directly performing second-round fine-tuning on the domain-general LoRA module with the domain-specific dataset? Unfortunately, the answer is no. Since LLMs generate natural language text, ensembling their outputs can introduce semantic inconsistencies or ambiguities, while also increasing inference time and GPU memory usage. Meanwhile, performing a second round of fine-tuning risks catastrophic forgetting (Kirkpatrick et al., 2017), causing the model to collapse.
Mixture-of-LoRA. We propose a simple yet effective method called Mixture-of-LoRA, which linearly merges the model parameters of the domain-general LoRA module and the domain-specific LoRA module . Formally, given the domain-general LoRA module and the domain-specific LoRA module , we define the mixture-of-LoRA operator as:
where the coefficients and represents the importance of merging. We constraint and . They can be considered hyperparameters and selected using the validation data. Please note that the mixed maintains the same total number of parameters as one standard LoRA, making our Mixture-of-LoRA method simple, fast, and effective. There is no extra cost at inference time in terms of memory or compute, since we only do element-wise operations on model weights. In addition, the domain-general LoRA module is reusable. When facing a new domain, it is only necessary to retrain a domain-specific LoRA module.
Mixture-of-LoRA is inspired by recent studies (Ilharco et al., 2023; Wortsman et al., 2022) on the linear connectivity of trained models in a full finetuning setting. These studies suggest that parameters of tuned models can be directly added to improve generalization, provided they are initialized from the same pre-trained model checkpoint. Specifically, as shown in Figure 3, recent research (Ilharco et al., 2023) defines the concept of ”task vector”. A task vector specifies a direction in the weight space of a pre-trained model , such that movement in that direction improves performance on the task. It is built by subtracting the weights of a pre-trained model () from the weights of the same model after fine-tuning () on a task. Adding task vectors together can improve performance on multiple tasks at once. The underlying hypothesis is that two models finetuned from the same pre-trained checkpoint often lie in the same error basin (Neyshabur et al., 2020; Zhang et al., 2023), and thus the parameters could be directly added. Extending this property to the context of LoRA, we hypothesize that LoRA modules can also be linearly combined. This is because a LoRA module can be considered as the difference between a fine-tuned LLM and its pre-trained counterpart, making it analogous to a task vector.
Entropy-Guided Adaptive Mixture-of-LoRA. As shown in Figure 2(c), in this subsection, we provide an efficient and automatic way to better choose mixture coefficients and . As discussed in the previous subsection, and can be chosen by employing the grid-search in the validation data. Nevertheless, (1) it is still lacking a guiding principle. (2) When the distribution of the inference data differs significantly from that of the validation set, the chosen coefficients may perform poorly. To this end, we introduce entropy minimization on the unlabeled test samples an optimization surrogate objective to update and . Specifically, the Shannon entropy (Shannon, 1948) is a well-known measure of uncertainty. For a sample , the predicted output of a neural network is , the Shannon entropy is calculated as , where denotes the probability that the input is predicted to be the -th class. Lower entropy indicates that the model has lower uncertainty about its predictions, meaning the model is more confident in its outputs. Therefore, the intuition behind our method is that the good coefficients and for the test inputs should make the mixed model more confident in its prediction, that is, it should lead to lower model entropy over the input (Wang et al., 2021a, b; Yang et al., 2024b). Formally, we collect a set of unlabeled test samples , i.e., some instruction inputs in the test time. We fix the , , , and using the following entropy minimization loss to update coefficients and :
For the LLM, the output of is a sentence. Since our instruction is to select a title from a given candidate set, the first few tokens output by the model are more important because after deciding on them, the subsequent tokens are more certain. So in practice, we only calculate the average entropy of the first three tokens in the sentence to represent . Besides, we do not need all test data to be available. Even if only 50 unlabeled tests are available, our method can have significant performance improvements.
3.4. Discussion
Key Advantages of MoLoRec. 1) Generalization. MoLoRec is generalizable to various recommendation scenarios. For example, when facing a new recommendation domain, it only needs to train a new domain-specific LoRA module, enabling rapid generalization. Even in extreme cases where no training data is available for the new domain, MoLoRec can still work using the generalizable base model. For the new user or new item recommendation scenario, MoLoRec dynamically balances generalization and domain-specific specialization to deliver accurate recommendations. Additionally, leveraging the in-context learning capabilities of LLMs, MoLoRec naturally exhibits task generalization. E.g., it can generate explainable recommendation results. 2) Efficiency. The proposed MoLoRec paradigm is designed for ease of use, offering plug-and-play integration, where the domain-general module is trained once, and the domain-specific plugin is incorporated with minimal fine-tuning.
Comparision to Existing Methods. Traditional sequential recommendation models (e.g., GRU4Rec (Hidasi et al., 2016), SASRec (Kang and McAuley, 2018)) typically rely on explicit item IDs for modeling, which restricts their generalization ability across new domains or platforms. To this end, transferable methods have been explored for universal sequence representation learning. For instance, VQ-Rec (Hou et al., 2023)) and UniSRec (Hou et al., 2022) leverage text to represent items and employ contrastive pre-training strategies on language models to enhance transferability. However, these pretraining methods cannot be directly applied to LLMs. With the rise of LLMs, their strong domain and task generalization capabilities present new opportunities for improving the generalization of recommender systems. We categorize these approaches as breadth-oriented methods. For example, P5 (Geng et al., 2022) designs prompts to unify five recommendation tasks, establishing a unified text-to-text recommendation paradigm. In contrast, depth-oriented methods focus on deeply aligning LLMs with specific domains. For instance, TallRec (Bao et al., 2023) employs LoRA for efficient domain adaptation. Another research direction explores aligning collaborative signals with LLMs, as seen in works like LLaRA (Liao et al., 2024) and iLoRA (Kong et al., 2024). While incorporating collaborative signals significantly enhances performance in warm-start scenarios, it comes at the cost of reduced generalization, making adaptation to new domains and cold-start situations more challenging.
4. Experiments
4.1. Experimental Settings
Methods Beauty Toys Sports MovieLens-1M NDCG@1 NDCG@3 NDCG@1 NDCG@3 NDCG@1 NDCG@3 NDCG@1 NDCG@3 Traditional BPR-MF 0.1630 0.2588 0.1276 0.2056 0.1496 0.2338 0.1724 0.4185 GRU4Rec 0.1672 0.2752 0.1320 0.2243 0.1787 0.2829 0.1724 0.4423 SASRec 0.2410 0.3284 0.2223 0.3105 0.1957 0.2967 0.2257 0.4708 FMLP-Rec 0.2988 0.4000 0.2994 0.3990 0.2645 0.3812 0.2410 0.5515 Transferable UniSRec 0.2654 0.4089 0.2612 0.3998 0.2341 0.3721 0.2615 0.5594 VQ-Rec 0.2714 0.4157 0.2715 0.4119 0.2476 0.3944 0.2805 0.5745 LLM-Based Qwen2-7B 0.0300 0.0394 0.0843 0.1062 0.0170 0.0242 0.0814 0.1057 RecFormer 0.2858 0.3840 0.3001 0.3880 0.2667 0.3885 0.2743 0.5701 P5 0.1775 0.2482 0.1171 0.1709 0.1860 0.2674 0.2046 0.2947 TALLRec 0.3208 0.3479 0.3308 0.3583 0.3002 0.3274 0.4759 0.4971 Ours MoLoRec-G 0.3081 0.3316 0.2957 0.3209 0.2750 0.2998 0.5680 0.5918 MoLoRec-S 0.4079 0.4291 0.4076 0.4314 0.3735 0.3925 0.5460 0.5703 MoLoRec 0.4132* 0.4350* 0.4097* 0.4334* 0.3754* 0.3944* 0.5783* 0.6023*
Methods Beauty Toys Sports MovieLens-1M NDCG@1 NDCG@3 NDCG@1 NDCG@3 NDCG@1 NDCG@3 NDCG@1 NDCG@3 Traditional BPR-MF 0.0306 0.0688 0.0333 0.0765 0.0350 0.0739 0.0723 0.1421 GRU4Rec 0.0562 0.1063 0.0447 0.0926 0.0640 0.0996 0.0798 0.1489 SASRec 0.0656 0.1368 0.0670 0.1210 0.0547 0.1203 0.0912 0.1891 FMLP-Rec 0.0587 0.1229 0.0537 0.1117 0.0545 0.1236 0.1145 0.1947 Transferable UniSRec 0.0957 0.1457 0.0814 0.1559 0.0832 0.1408 0.0985 0.1343 VQ-Rec 0.1189 0.1589 0.0957 0.1603 0.0985 0.1463 0.1025 0.1412 LLM-Based Qwen2-7B 0.0187 0.0260 0.0293 0.0356 0.0213 0.0273 0.0318 0.0407 RecFormer 0.1051 0.1687 0.0913 0.1592 0.0922 0.1489 0.1108 0.1547 P5 0.0871 0.1466 0.0755 0.1358 0.0758 0.1355 0.0957 0.1319 TALLRec 0.1415 0.1674 0.1251 0.1524 0.1226 0.1486 0.1458 0.1668 Ours MoLoRec-G 0.1746 0.2072 0.1474 0.1710 0.1581 0.1868 0.1455 0.1890 MoLoRec-S 0.1603 0.1863 0.1504 0.1764 0.1564 0.1867 0.1636 0.2597 MoLoRec 0.1825* 0.2145* 0.1614* 0.1821* 0.1646* 0.1921* 0.1818* 0.2755*
4.1.1. Datasets.
We conduct experiments on e-commerce and movie recommendation scenarios. For the e-commerce recommendation scenario, the domain-general instruction tuning dataset is conducted using seven e-commerce domains in Amazon111https://jmcauley.ucsd.edu/data/amazon/. and validated on three domain-specific datasets in Amazon (Beauty, Toys, Sports). For the movie recommendation scenario, the domain-general dataset is built using MovieLens-10M222https://grouplens.org/datasets/movielens/ and validated on the domain-specific dataset MovieLens-1M.
For all datasets, items are represented using their textual ”title” information. To prevent data leakage, we carefully removed the overlapping portions between the domain-general dataset and the domain-specific datasets. We consider two recommendation settings: 1) Warm-Start Setting keeps the five-core dataset and filters users and items with fewer than five interactions for all datasets. Following (Geng et al., 2022; Lin et al., 2024b), we adopt the leave-one-out strategy to split the filtered dataset. More concretely, we split the last interaction of each user into the test set, the second-to-last one into the validation set, and the rest into the training data. 2) New-Item Setting uses the same training and validation sets as the warm-start setting, but replaces the items in the test set with those that never appear in the training or validation sets. (See Appendix A.1 for dataset statistics.)
4.1.2. Baselines.
We compare MoLoRec with traditional recommendation methods (BPR-MF (Rendle et al., 2009), GRU4Rec (Hidasi et al., 2016), SASRec (Kang and McAuley, 2018), and FMLP-Rec) (Zhou et al., 2022), transferable sequential recommenders (UniSRec (Hou et al., 2022), VQ-Rec (Hou et al., 2023)), LLM-based recommenders (Qwen2-7B (Yang et al., 2024c), RecFormer (Li et al., 2023), P5 (Geng et al., 2022), TALLRec (Bao et al., 2023)) and our ablation counterparts (MoLoRec-G, MoLoRec-S). (See Appendix A.2 for more details of these baselines.)
4.1.3. Evaluation Setting.
Following some previous LLM-based recommendation works (Zhang et al., 2024; Kim et al., 2024b), to evaluate the performance of the sequential recommendation models, we add 29 randomly selected non-interacted items to the test set, so that the test set of each user contains 1 positive item and 29 negative items. For quantitative comparison, we employ widely used ranking-based metrics, NDCG@1 and NDCG@3 for all experiments. All metrics are ”the higher, the better”. For all tables in the following, bold* numbers refer to the best performance, while underlined numbers indicate the second-best performance.
(See Appendix A.3 for implementation details.)
4.2. Overall Performance
We comprehensively compare MoLoRec against traditional, transferable, and LLM-based recommenders. The experimental results on the warm-start I.I.D setting and the cold-start item out-of-distribution settings are shown in Table 1 and 2, respectively. From the experimental results, we have the following observations:
The proposed MoLoRec consistently achieves the best performance across all I.I.D and O.O.D scenarios on the four datasets, with a t-test at p¡0.05 level. Specifically, in the warm-start scenario, MoLoRec achieves notable improvements in NDCG@1 over the best baseline methods (excluding our ablation counterparts), with performance gains of 28.8%, 23.85%, 25.05%, and 21.52% across the Beauty, Toys, Sports, and Movielens-1M, respectively. In the cold-start O.O.D scenario, the improvements are 28.98%, 29.02%, 34.26%, and 24.69%. The outstanding performance of MoLoRec in both I.I.D and O.O.D scenarios demonstrates its ability to efficiently capture domain-specific knowledge while exhibiting exceptional generalization capabilities.
Qwen2-7B demonstrates limited performance across all scenarios. However, the LLMs trained via TALLRec achieved significant improvements. This is because there is a gap between pre-training general text corpus of LLMs and the recommendation task, showing the importance of using recommendation knowledge for instruction fine-tuning on pre-trained LLMs.
Traditional recommendation methods and the ID-based LLM recommendation method P5 perform poorly in cold-start scenarios. Relying heavily on collaborative filtering information reduces the model’s generalization capability.
4.3. In-Depth Analysis
4.3.1. Ablation Study
From the performance of MoLoRec, MoLoRec-G, MoLoRec-S in the Table 1 and 2, we can find that: 1) MoLoRec consistently surpasses its ablation counterparts in all scenarios, which indicates that it is crucial for LLM to comprehend both general recommendation world knowledge and domain-specific knowledge. These two types of knowledge can complement each other. It also validates that the mixture-of-LoRA method can effectively integrate both types of knowledge. 2) Note that the MoLoRec-G has not been exposed to training data from the Beauty, Toys, Sports, or Movielens-1M domains in any experimental setting. However, it still achieves commendable performance in such zero-shot settings, demonstrating that it has effectively learned generalizable knowledge in the recommendation domain. This strong performance underscores its robust generalization ability.
4.3.2. Performance in Few-Shot Training Setting
We further conduct experiments in scenarios with limited domain-specific training data. Specifically, we adopt a few-shot training setup on MovieLens-1M, where only a small percentage of samples are randomly selected from the training set for model training. We compare MoLoRec with TallRec and the results of second-round fine-tuning on the generalizable base model. The experimental results are shown in Table 3. We find that the optimization approach of second-round fine-tuning led to catastrophic forgetting. It fails to generate output in the specified instruction format. Experimental results demonstrate that MoloRec maintains strong performance even in few-shot scenarios.
4.3.3. Analysis of the Coefficients and
In Figure 4(a) and 4(b), we investigate the coefficients and calculated by entropy-guided adaptive mixture-of-LoRA in the warm-start scenario and the cold-start scenario, respectively. and represent the respective weights assigned to the domain-general and domain-specific LoRA module during their mixture. We observe that in the warm-start scenario, is relatively large, reflecting a greater reliance on domain-specific LoRA. Conversely, in the cold-start scenario, the weight of increases significantly, emphasizing the importance of domain-general knowledge. This result is reasonable and aligns with the differing requirements of these two scenarios. This result also demonstrates the effectiveness of our entropy-guided adaptive mixture-of-LoRA method.
4.3.4. Analysis of the Number of Unlabeled Test Data
The number of unlabeled test data is one of the hyperparameters. Figure 5 illustrates the impact of different numbers on the NDCG@1 performance. The entropy-guided learning method converges rapidly, experimental results indicate that setting the number to 50 or 100 achieves a model fusion weight with optimal performance, this configuration proves effective across the majority of experiments conducted on the Beauty, Toys, Sports and MovieLens-1M datasets. Therefore, a limited amount of training data is sufficient to learn suitable model fusion weights, thereby significantly reducing the data requirements and computation resource costs.
4.3.5. Case Study
We further conduct a case study to delve deeper into the recommendation results of MoLoRec. We randomly selected a user from the Movielens-1M test set, provided their historical viewing records and a candidate set, and asked both our model and ChatGPT to make movie recommendations along with explanations for their choices. The outputs are shown in Figure 6. We find that MoLoRec successfully generalizes to the explainable recommendation task. It accurately captured the user’s preference for action movies from their historical viewing records and leveraged world knowledge to provide an accurate interpretation of the plot of Die Hard. In contrast, GPT-4 lacks domain-specific knowledge in the recommendation, incorrectly associating the action movie Die Hard with science fiction films like Star Wars, Alien, and Terminator, resulting in unreasonable explanation outcomes. This case study further highlights MoLoRec’s task generalization capability and its deep understanding of recommendation knowledge.
Scenario | Sample % | TallRec | 2nd Finetune | MoLoRec |
Warm-Start | 10% | 0.3957 | 0.0704 | 0.5353 |
20% | 0.4298 | 0.0790 | 0.5454 | |
30% | 0.4563 | 0.0540 | 0.5498 | |
Cold-Start | 10% | 0.1091 | 0.0000 | 0.1455 |
20% | 0.1091 | 0.0182 | 0.1636 | |
30% | 0.1273 | 0.0000 | 0.1636 |
5. Related Works
Sequential Recommendation. Sequential recommendation utilizes users’ interaction histories to predict the next relevant item. Deep learning-based methods, such as RNNs (Hidasi et al., 2016; Li et al., 2017), GNNs (Qiu et al., 2019; Chang et al., 2021), and attention mechanisms (Kang and McAuley, 2018; Zhou et al., 2020; Xie et al., 2022), have become mainstream but rely solely on item IDs, limiting their adaptability to new scenarios. To improve transferability, transferable sequential recommendation studies (Hou et al., 2022, 2023; Li et al., 2023; Yuan et al., 2023) explore leveraging textual features to enhance item representations. These approaches improve the transferability and robustness of recommender systems. Nowadays, LLMs offer new opportunities for sequential modeling, promising more robust and generalizable recommender systems.
LLM-Based Recommendation. With the rise of LLMs, interest in LLM-based recommender systems has grown, leveraging LLMs as core engines. Early studies (Dai et al., 2023; Sanner et al., 2023; Wang et al., 2023) explore their zero-shot/few-shot potential via in-context learning (Dong et al., 2024). However, the gap between LLMs’ pretraining on general text and recommendation-specific needs leads to suboptimal performance. To address this, recent research follows two paradigms: breadth-oriented and depth-oriented. The former integrates multi-domain (Tang et al., 2024; Peng et al., 2024) or multi-task (Geng et al., 2022; Zhang et al., 2024; Cui et al., 2022; Peng et al., 2024) recommendation data to construct extensive recommendation world knowledge, paving the way for developing a generalizable LLM-based recommender. For example, Peng et al.(Peng et al., 2024) build large-scale e-commerce instruction dataset ECInstruct and develop generalist LLM for e-commerce. P5 (Geng et al., 2022) designs prompts to unify 5 recommendation tasks and presents a unified text-to-text recommendation paradigm. Depth-oriented paradigm seeks to enable LLMs to deeply comprehend recommendation tasks within specific domains. Key areas of focus include: 1) the in-depth alignment of domain-specific recommendation knowledge, such as collaborative signals (Lin et al., 2024b; Kim et al., 2024a; Liao et al., 2024; Ren et al., 2024; Kong et al., 2024). The introduction of collaborative signals effectively improves model performance in warm-start scenarios. However, this comes at the cost of reduced generalization, making it challenging to adapt to new domains and cold-start situations. Another research focus is 2) the development of efficient alignment methods between large language models and recommendation tasks. These methods include leveraging (Bao et al., 2023; Kong et al., 2024) LoRA fine-tuning techniques and designing data-efficient fine-tuning strategies (Lin et al., 2024c; Zheng et al., 2024). In fact, the two paradigms have complementary advantages. In this work, we investigate how to integrate the advantages of both paradigms to simultaneously achieve both breadth and depth.
Model Merging. Model merging aims to combine multiple expert models into a more capable single model, offering benefits such as reduced storage and serving costs (Yang et al., 2024a). Model merging techniques are applied in various scenarios such as unlearning old-knowledges in LLMs (Zaman et al., 2024), understanding content across multiple modalities (Aiello et al., 2024; Chen et al., 2024), generating images with different styles or achieving image-style transformation (Biggs et al., 2024). Previous attempts involve merging multiple models, all initially trained on the same task, with the aim of enhancing the model’s overall generalization (Gupta et al., 2020; Wang et al., 2022). Inspired by these works, we apply model merging to the LoRA modules, leveraging their ability to integrate domain-general and domain-specific knowledge effectively.
6. Conclusion
In this paper, we proposed a generalizable and efficient LLM-based recommendation framework MoLoRec. As the recommendation data for a specific domain is limited, MoLoRec is designed to combine domain-general recommendation knowledge and domain-specific recommendation knowledge. Specifically, MoLoRec first aligned LLM with the general recommendation task by fine-tuning LLM from a multi-domain instruction dataset, and then modeled domain-specific recommendation tasks by fine-tuning the domain-specific instruction dataset. After that, the combination is achieved by merging two fine-tuning model parameters. MoLoRec is generalizable as it injects recommendation general knowledge to any domain-specific recommendation tasks. Besides, MoLoRec is efficient not only because it chooses parameter-efficient fine-tuning, but also the plug-and-play nature of any domain-specific recommendation task and domain-general task. Extensive experimental results on four recommendation datasets under both the warm scenario and cold-start scenario show the effectiveness of our proposed framework. E.g., MoLoRec shows about 30% average improvement on four datasets under cold-start scenarios.
Appendix A Appendix: Experimental Details
Domain-General Dataset-1 | ||||
Datasets | # Users | # Items | # Interactions | Density(%) |
Clothing | 39,387 | 23,033 | 278,677 | 0.0307 |
Cell | 27,879 | 10,429 | 194,439 | 0.0669 |
Grocery | 14,681 | 8,713 | 151,254 | 0.1182 |
Health | 38,609 | 18,534 | 346,355 | 0.0484 |
Home | 66,519 | 28,237 | 551,682 | 0.0294 |
Pet | 19,856 | 8,510 | 157,836 | 0.0934 |
Tools | 16,638 | 10,217 | 134,476 | 0.0791 |
Videos | 24,303 | 10,672 | 231,780 | 0.0894 |
Domain-Specific Dataset-1 | ||||
Beauty | 22,363 | 12,101 | 198,502 | 0.0734 |
Toys | 19,412 | 11,924 | 167,597 | 0.0724 |
Sports | 35,598 | 18,357 | 296,337 | 0.0453 |
Domain-General Dataset-2 | ||||
Movielens-10M | 71,567 | 10,681 | 10,000,054 | 1.3082 |
Domain-Specific Dataset-2 | ||||
Movielens-1M | 6,040 | 6,883 | 1,000,209 | 2.4059 |
A.1. Dataset
The statistics of the domain-general and the domain-specific datasets are shown in Table 4.
A.2. Baselines
We compare MoLoRec with traditional recommendation methods (BPR-MF, GRU4Rec, SASRec, and FMLP-Rec), transferable sequential recommenders (UniSRec, VQ-Rec), LLM-based recommenders (Qwen2-7B, RecFormer, P5, TALLRec) and our ablation counterparts (MoLoRec-G, MoLoRec-S).
BPR-MF (Rendle et al., 2009) is one of the most representative collaborative filtering models.
GRU4Rec (Hidasi et al., 2016) is a seminal method that uses RNNs to model user action sequences for session-based recommendation.
SASRec (Kang and McAuley, 2018) is a representative sequential recommender model that adopts a self-attention mechanism to learn the item dependency from user interactions.
FMLP-Rec (Zhou et al., 2022) is an all-MLP model with learnable filters for sequential recommendation tasks.
UniSRec (Hou et al., 2022) equips textual item representations with an MoE-enhanced adaptor for domain fusion and adaptation. Both item-sequence and sequence-sequence contrastive learning tasks are designed for pre-training transferable sequence representations.
VQ-Rec (Hou et al., 2023) learns vector-quantized item representations for transferable sequential Recommenders.
Qwen2-7B333https://huggingface.co/Qwen/Qwen2-7B-Instruct is a well-known open-source LLM. In our experiments, we choose it as MoLoRec ’s LLM backbone.
P5 (Geng et al., 2022) is a unified LLM-based recommendation framework. It is built on T5 by fine-tuning with multiple recommendation tasks.
TALLRec (Bao et al., 2023) learns the recommendation task based on prompts consisting solely of text and fine-tunes the LLMs using the LoRA.
MoLoRec-G is an ablation counterpart of our proposed framework. It only underwent stage 1, utilizing only the domain-general LoRA module.
MoLoRec-S is an ablation counterpart of our proposed framework. It only underwent stage 2, utilizing only the domain-general LoRA module.
A.3. Implementation Details
To ensure a fair comparison, the experimental settings are standardized as follows: For traditional recommendation methods (BPR-MF, GRU4Rec, SASRec, and FMLP-Rec), the learning rate is set to 0.001, and the Adam optimizer is employed. The batch size is set to 256, and the embedding dimension is set to 64. Regarding transferable sequential recommenders (UniSRec, VQ-Rec), these models utilize a BERT for text processing. Specifically, the pre-trained models provided by the original authors are fine-tuned on our dataset. For RecFormer, the pre-trained model provided by the original work is also fine-tuned on the downstream tasks. In the case of P5 and TALLRec, an identical instruction fine-tuning template is used to align the original models with the recommendation tasks. Here, P5 undergoes full fine-tuning, while TALLRec is fine-tuned using a LoRA approach with a rank of 16. This setup standardizes the evaluation framework across different recommendation methodologies, ensuring comparability and fairness in assessing their performance. We use Qwen2-7B as the LLM backbone for MoLoRec. For parameter-efficient finetuning(PEFT) conducted on NVIDIA RTX 4090(24G) GPUs, we adopt low-rank adaption(LoRA) with LoRA rank as 16, LoRA alpha as 32, and LoRA dropout as 0.05 to get general LoRA adapter and target LoRA adapter. The learning rate is selected from 1e-4,2e-4 and the batch size is set to 128. For LoRA adapters fusion weights learning conducted on two NVIDIA RTX 4090(24G) GPUs, the batch size is set to 60, number of test samples for training is selected from 50,100. The number of tokens at the beginning of each title involved in training is set to 3 for Toys and Sports and 5 for Beauty because titles in Beauty are longer than Toys and Sports. In order to reduce GPU memory usage, we employed gradient checkpointing techniques. Since the model structure has not changed, we use the VLLM inference acceleration framework to perform inference and then evaluate the results. Other implementation details are available in our open-source code.